OpenStudy (anonymous):

How many solutions does this system have? y = x + 5 y = –5x – 1 one none infinite two

jimthompson5910 (jim_thompson5910):

y = x + 5 -5x-1 = x + 5 ... plug in y = -5x - 1 now solve for x

OpenStudy (anonymous):

2nd method is..plot the graph of both functions (straight lines ) intersection points will be the solution .. You can solve any qu. of this type by plotting graph. ( clearly two straight lines can intersect only at one point so your ans . is one).

OpenStudy (anonymous):

x + 5 = -x - 1 +1 +1 x + 6 = -x -x -x (?) 6 = -2 x?

OpenStudy (anonymous):

or -x = 3?

jimthompson5910 (jim_thompson5910):

well it's not x = -3, it should be x = -1 so this suggests that there is only one solution